OpenStudy (abmon98):

A ball moves on the horizontal surface of a billiards table with deceleration of constant magnitude d m s−2 . The ball starts at A with speed 1.4 m s−1 and reaches the edge of the table at B, 1.2 s later, with speed 1.1 m s−1 . (i) Find the distance AB and the value of d. [3] AB is at right angles to the edge of the table containing B. The table has a low wall along each of its edges and the ball rebounds from the wall at B and moves directly towards A. The ball comes to rest at C where the distance BC is 2 m.

OpenStudy (abmon98):

(ii) Find the speed with which the ball starts to move towards A and the time taken for the ball to travel from B to C. [3] (iii) Sketch a velocity-time graph for the motion of the ball, from the time the ball leaves A until it comes to rest at C, showing on the axes the values of the velocity and the time when the ball is at A, at B and at C

OpenStudy (anonymous):

(i) \(v_f=v_i+at\). \(v_f\) is the final velocity, \(v_i\) is the initial velocity, \(a\) is the acceleration\deceleration, \(t\) is the period of time. \(1.1=1.4+a\times 1.2 \rightarrow a = -0.25 m/s^2\) \(v_f^2-v_i^2=2as\). \(s\) is the distance = AB \(\rightarrow 1.1^2-1.4^2=2\times -0.25 \times s \rightarrow s =1.5 m\)

OpenStudy (anonymous):

(ii) Is the deceleration when the ball moves from B to C \(-0.25m/s^2\)?

OpenStudy (abmon98):

i solved i) the problem is with understanding ii) a bit

OpenStudy (anonymous):

Great. Just a little effort and we can do everything.

OpenStudy (abmon98):

i want to know why the acceleration to approach the wall is equal when the ball rebounds

OpenStudy (anonymous):

I don't quite sure. I just searched for the answer but couldn't find an appropriate one for this.

OpenStudy (abmon98):

i solved the question

OpenStudy (abmon98):

but i dont understand why the acceleration before is the same acceleration when rebounded.

OpenStudy (abmon98):

the deceleration to approach B is the same deceleration after re bouncing. -1/2(-.25)t^2=2 0.125t^2=2 t=4seconds a=v-u/t-->-0.25*4=0-u u=1m/s

OpenStudy (anonymous):

Yes, you are right.

OpenStudy (anonymous):

I had the same question when I studied about this. Why the acceleration is the same???

OpenStudy (abmon98):

@ProfBrainstorm